# protoplanetary disks — planets and satellites: composition

### v1

@@ 2.1 Overview
Figure 1: Schematic illustration of our model. We consider a protoplanetary disk consisting of gas and dust. The dust grains contain ammonium salts and NH3 ice as nitrogen carriers. The grains grow through mutual collisions and drift inward. When the grains crosses the sublimation/dissociation line, they release NH3 vapor. A protoplanet forms at given time and orbit. This protoplanet contains ammonium salts in its core, and core erosion releases NH3 into its atmosphere. The planet also accretes NH3 vapor-rich disk gas during the gas accretion phase.
@@ 2.2 Disk model
$\frac{\partial\Sigma_{\rm g}}{\partial t}=-\frac{1}{r}\frac{\partial}{\partial r}(rv_{\rm g}\Sigma_{g}),$ (1)
@@
where $M_{\rm disk}$ is the total mass of the disk and $r_{\rm c}$ is the characteristic radius of the disk. We assume a compact disk with $M_{\rm disk}=0.05M_{\odot}$ and $r_{\rm c}=50~{}\rm{au}$.
@@ 2.3 Dust evolution
$\frac{\partial\Sigma_{\rm d}}{\partial t}+\nabla\cdot\mathcal{M_{\rm d}}=-\sum_{i}S_{\rm{N},i},$ (7)
@@
where $v_{\rm d}$ and $D_{\rm d}$ are the radial velocity and diffusion coefficient of the dust, respectively, and $S_{\rm{N},i}$ represents the rate of dust surface density loss due to ammonium salt dissociation or NH3 ice sublimation (the subscript $i$ stands for either ammonium salt or NH3 ice). We describe our prescription for $S_{\rm{N},i}$ in Section 2.4. The dust radial velocity is given by (Whipple, 1972; Adachi et al., 1976; Weidenschilling, 1977)
@@ 2.4 Production and transport of nitrogen-bearing vapor
$\frac{\partial N_{i,{\rm vap}}}{\partial t}+\nabla\cdot\mathcal{N}_{i,{\rm vap}}=\sum_{i}S_{\rm{N},i},$ (19)
@@
In this study, we adopt the following simplified prescription for $S_{\rm{N},i}$,
$S_{\rm{N},i}=\left\{\begin{array}[]{ll}{\displaystyle f_{i,\rm{dust}}}{\displaystyle\frac{\mathcal{M}_{\rm{d}}}{m_{\rm i}r\Delta r}},&T=T_{\mathrm{vap},i},\\ 0,&\rm{otherwise},\end{array}\right.$ (21)
where $f_{i,\rm{dust}}$, $m_{\rm i}$, and $T_{\mathrm{vap},i}$ are the mass fraction in the dust, the molecular mass, and dissociation/sublimation temperature of the nigrogen-bearing species $i$, respectively, and $\Delta r$ is the cell width. We assume that sublimation or dissociation occurs immediately when the dust grains cross the NH3 ice line or the ammonium salt line, releasing NH3 vapor. NH3 vapor can be converted to highly volatile chemical species such as atomic N and N2 by photodissociation (Schwarz & Bergin, 2014). The disk observations of Pontoppidan et al. (2019) support that NH3 depletion in the inner region of the disk ($r<10$ au) comes from the destruction of NH3 vapor by photodissociation. We neglect recondensation of the produced NH3 vapor outside the NH3 ice line in a fiducial model.
@@ 2.5 Planetary nitrogen abundance
We calculate the planetary nitrogen abundance sequentially from the disk evolution. For computational simplicity, we insert a protoplanet at time $t=t_{\rm p}$ and orbital radius $r=r_{\rm p}$, reset the inner boundary of the gas disk to the core’s orbit, and let the core accrete all the disk gas flowing to the inner computational boundary by applying the zero-torque boundary condition. The zero-torque boundary condition of the gas produces a pressure bump outside the inner boundary, which causes dust pile-up. This effect reproduces gap formation. We neglect core growth, planetary migration, and gas inflow from the inner region of the core’s orbit. This treatment is based on the runaway accretion scheme of Tanigawa & Tanaka (2016); Tanaka et al. (2020). They show analytically that a planet core with a 30 $M_{\oplus}$$10M_{\rm jup}$ would capture almost all gas inflowing into the planetary gap, where $M_{\rm jup}$ represents Jupiter’s mass. In our calculations, the protoplanet is assumed to have the initial mass $M_{\rm p,ini}$ of 30 $M_{\oplus}$ and comprises a heavy element core and an envelope. The mass of the heavy-element core is taken to be 13 $M_{\oplus}$ (0.04 $M_{\rm jup}$) based on Juno’s constraint that sets an upper limit of 6–27 $M_{\oplus}$ for the total heavy-element mass of Jupiter (Wahl et al., 2017; Ni, 2019). However, Jupiter’s internal structure model besed on the Juno’s gravity data suggests that Jupiter’s core is diluted (e.g., Vazan et al., 2018; Debras & Chabrier, 2019; Helled et al., 2022). Furthermore, Jupiter’s formation models show that once the core grows into a few $M_{\oplus}$ in mass and acquires an envelop, heavy elements tend to dissolve in the envelope, irrespective of whether the solids are brought about by pebble accretion or planetesimal accretion (Iaroslavitz & Podolak, 2007; Hori & Ikoma, 2011; Brouwers et al., 2018; Valletta & Helled, 2020). Therefore, we assume that 10 $M_{\oplus}$ out of 13 $M_{\oplus}$ of heavy elements in the core dissolve into the atmosphere. Several long-term evolutionary models of Jupiter’s interior structure have demonstrated that convective mixing erases compositional gradients in the outer regions of the planet (50% or more of the radius) on scales of millions of years (Vazan et al., 2018; Müller et al., 2020). We assume that heavy elements dissolved in the atmosphere undergo homogenization through convective mixing. On the other hand, we note that the composition of actual Jupiter is not entirely uniform in the vertical direction, as evidenced by Jupiter having a diluted core. A discussion of compositional heterogeneity is provided in Section 4.6.
@@
We continue calculating the time evolution of dust and NH3 vapor using equations (7), (13) and (19), but with the inner boundary condition at $r=r_{\rm p}$. Because the ammonium salt line is located inside the planetary core orbit, salt dissociation does not contribute to NH3 vapor abundance after core formation. For the dust flux at the inner boundary, we adopt a zero-flux condition because of the lack of background gas inside the boundary, i.e., no radial drift. The inner boundary flux of NH3 vapor is determined solely from the advection term: the first term on the right-hand side of equation (20).
@@ 2.6 Numerical settings and parameter set
The simulation consists of several models with different parameter sets. We set up as a fiducial model when a core forms at 0.6 Myr at 3.0 au in a disk with moderate turbulence strength $\alpha=10^{-3}$, a fragmentation velocity of icy dust $v_{\rm frag}=10$ m s-1, and salt dissociation temperature $T_{\rm vap,NH_{4}X}=200$ corresponding to ${\rm X}={\rm HCO_{2}}$. Other parameter sets are shown in Table 1, with seven cases in which the protoplanet formation age, core orbit, salt dissociation temperature, turbulence strength, or dust fragmentation velocity differ from the fiducial model. In addition to these parameter sets, we calculate Case 8, which considers the effect of recondensation according to equations (22) and (23). For each case, we take the ammonium salt content in the dust to be 10, 20, or 30 wt(corresponding to $f_{\rm NH_{4}X,dust}=0.2^{+0.1}_{-0.1}$).
@@ 3.1.1 Time evolution of the disk until the core formation
Figure 2: Time evolution of the disk temperature $T$ from the fiducial model as a function of the distance from the central star $r$. The blue dotted, dashed, and solid lines are the snapshots at times $t=0,0.1$ and 0.6 Myr, respectively.
@@
Figure 3: Time evolution of the gas surface density $\Sigma_{\rm g}$ (upper left), the dust surface density $\Sigma_{\rm d}$ (upper right), the Stokes number St (lower left), and the dust grain size (lower right), respectively. The results are from the fiducial model. The blue dotted, dashed, and solid lines are the snapshots at times $t=0,0.1$ and 0.6 Myr, respectively. The small drops in $\Sigma_{\rm d}$ at 2.5 au (on the dashed line) and 1.8 au (on the solid line) are caused by ammonium salt dissociation.
@@
Figure 4: Time evolution of the number of NH3 per unit $\ln r$ derived from NH4HCO2 salt $N_{\rm NH_{4}HCO_{2},vap}^{\prime}$ (green) and NH3 ice $N_{\rm NH_{3},vap}^{\prime}$ (blue) from the fiducial model. The dashed and solid lines are the snapshots at $t=0.1$ and 0.6 Myr. The open and filled circles above the panel indicate the locations of the NH3 ice line (gray) and the ammonium salt line (black) at $t=0$ and $t=0.6$ Myr.
@@ 3.1.2 Gas accretion onto the protoplanet
Figure 5: Time evolution of the gas surface density $\Sigma_{\rm g}$ (left) and $N_{i,\rm vap}^{\prime}$ (right) after the core insertion. The core is inserted at $t_{\rm p}=$ 0.6 Myr and $r_{\rm p}=$ 3.0 au. These results are from the fiducial model. The dotted, dashed, and solid lines are the snapshots at times $t-t_{\rm p}=0,0.01$ and 0.04 Myr ($M_{\rm p}=M_{\rm jup}$), respectively. Here, $M_{\rm p}$ is the total planetary mass as $M_{\rm p}=M_{\rm p,atm}+M_{\rm p,core}$.
@@
Figure 6: Mass of the planet as a function of time after core insertion $t-t_{\rm p}$ (left) and atmospheric nitrogen abundance of the planet $x_{\rm{N,p}}$ normalized by solar value of $7.78\times 10^{-5}$ (Asplund et al., 2021) as a function of $M_{\rm p}$ (right) from the fiducial model. The orange line in the right panel shows the planetary nitrogen abundance $x_{\rm N,p}$, with the orange shaded area representing the 10 wtuncertainty in the NH4HCO2 salt content in comets (Poch et al., 2020). The blue line indicates the contribution of accreted NH3 vapor only ($n_{\rm NH_{3},atm}/(n_{\rm H,atm}+3n_{\rm NH_{3},atm})$). The black hatched bar marks Jupiter’s atmospheric nitrogen abundance including observational errors from Juno (Bolton et al., 2017). The dot-dashed line marks Jupiter’s nitrogen abundance from Galileo probe (Wong et al., 2004); this value is highly uncertain ($4.2^{+1.6}_{-1.6}$), but the uncertainty is not indicated in the figure for clarity. The light gray line marks the lower limit of Saturn’s nitrogen abundance from Cassini (Fletcher et al., 2009).
@@
Here we describe disk evolution after core insertion and gas accretion by the core. Figure 5 shows the time evolution of $\Sigma_{\rm g}$ and $N_{i,\rm vap}^{\prime}$ as a function of the time $t-t_{\rm p}$ after core insertion. The NH3 ice and ammonium salt lines are located outside and inside the core orbit, respectively. Although there is a small supply of NH3 vapor after the core insertion, the amount of nitrogen vapor that accretes on the planet is largely determined by the amount of vapor that has reached the outer core orbit prior to the core insertion. The left panel of Figure 6 displays the evolution of the planet mass. The planet mass reaches one Jupiter mass $M_{\rm jup}$ within 0.04 Myr after the onset of gas accretion.
@@
The right panel of Figure 6 shows the nitrogen enrichment of the planet $x_{\rm N,p}$ as a function of $M_{\rm p,atm}$. Here we plot the possible range of nitrogen abundance that was originally derived from the dissociation of the ammonium salt, sublimation of NH3 ice, and ammonium salt in the core, taking into account the 10 wtuncertainty in the abundance of salts contained in the disk dust. To highlight the contribution of NH3 vapor to nitrogen enrichment, we also plot the nitrogen abundance of accreting gas only. The final nitrogen abundance in the planetary atmosphere exceeds the solar abundance by a factor of 1–3. Dust containing 20 wtof ammonium salts yields $x_{\rm N,p}$ comparable to that measured by Juno for Jupiter (Bolton et al., 2017). Salts incorporated into the core and accreted NH3 vapor almost equally contribute to this degree of nitrogen enrichment; either one alone results in the enrichment degree below Juno’s observations.

### v2

# Nitrogen transport in protoplanetary disks by ammonium salts: a possible origin of Jupiter’s nitrogen enrichment

@@ 2.1 Overview
Figure 1: Schematic illustration of our model. We consider a protoplanetary disk consisting of gas and dust. The dust grains contain ammonium salts and NH3 ice as nitrogen carriers. The grains grow through mutual collisions and drift inward. When the grains crosses the sublimation/dissociation line, they release NH3 vapor. A protoplanet forms at given time and orbit. This protoplanet contains ammonium salts in its core, and core erosion releases NH3 into its atmosphere. The planet also accretes NH3 vapor-rich disk gas during the gas accretion phase. Alt text: Color illustration of our model.
@@ 2.2 Disk model
$\frac{\partial\Sigma_{\rm g}}{\partial t}=-\frac{1}{r}\frac{\partial}{\partial r}(rv_{\rm g}\Sigma_{\rm g}),$ (1)
@@
where $M_{\rm disk}$ is the total mass of the disk and $r_{\rm c}$ is the characteristic radius of the disk. We assume a compact disk with $M_{\rm disk}=0.05M_{\odot}$ and $r_{\rm c}=50~{}\rm{au}$. This value represents a typical disk mass and is also consistent with the gas mass estimated by the minimum-mass solar nebula model within a factor of three (Hayashi, 1981).
@@ 2.3 Dust evolution
$\frac{\partial\Sigma_{\rm d}}{\partial t}+\nabla\cdot\mathcal{M_{\rm d}}=-\sum_{i}S_{\mathrm{N},i},$ (7)
@@
where $v_{\rm d}$ and $D_{\rm d}$ are the radial velocity and diffusion coefficient of the dust, respectively, and $S_{\mathrm{N},i}$ represents the rate of dust surface density loss due to ammonium salt dissociation or NH3 ice sublimation (the subscript $i$ stands for either ammonium salt or NH3 ice). We describe our prescription for $S_{\mathrm{N},i}$ in Section 2.4. The dust radial velocity is given by (Whipple, 1972; Adachi et al., 1976; Weidenschilling, 1977)
@@
By solving equations (7) and (13), we calculate the dust surface density and dust size using $m_{\rm d}=\Sigma_{\rm d}/N_{\rm d}$ and $m_{\rm d}=4\pi a^{3}\rho_{\rm s}/3$.
@@ 2.4 Production and transport of nitrogen-bearing vapor
$\frac{\partial N_{i,{\rm vap}}}{\partial t}+\nabla\cdot\mathcal{N}_{i,{\rm vap}}=\sum_{i}\left(\frac{S_{\mathrm{N},i}}{m_{i}}\right),$ (19)
@@
where $m_{\rm i}$ is the molecular mass of the nigrogen-bearing species $i$.
In this study, we adopt the following simplified prescription for $S_{\mathrm{N},i}$,
$S_{\mathrm{N},i}=\left\{\begin{array}[]{ll}{\displaystyle f_{i,\rm{dust}}}{\displaystyle\frac{\mathcal{M}_{\rm{d}}}{r\Delta r}},&T=T_{\mathrm{vap},i},\\ 0,&\rm{otherwise},\end{array}\right.$ (21)
where $f_{i,\rm{dust}}$, and $T_{\mathrm{vap},i}$ are the mass fraction in the dust and dissociation/sublimation temperature of the nigrogen-bearing species $i$, respectively, and $\Delta r$ is the cell width. This formulation implies that solid nitrogen molecules contained in dust are instantaneously converted into vapor upon crossing the NH3 ice line or the ammonium salt line. In general, similar models that describe the evolution of dust composition determine sublimation and recondensation rates based on the saturation vapor pressure of each chemical species (Schoonenberg & Ormel, 2017; Booth & Ilee, 2019; Schneider & Bitsch, 2021). Our model corresponds to a case where the recondensation rate is zero at the snow and dissociation lines, while the sublimation rate is sufficiently high for the chemical species to be immediately lost from the dust. Vapor release occurs only at the NH3 ice line or at the ammonium salt line. Furthermore, we assume that NH3 and ammonium salts are absent within these orbits under initial conditions. NH3 vapor can be converted to highly volatile chemical species such as atomic N and N2 by photodissociation (Schwarz & Bergin, 2014). The disk observations of Pontoppidan et al. (2019) support that NH3 depletion in the inner region of the disk ($r<10$ au) comes from the destruction of NH3 vapor by photodissociation. We neglect recondensation of the produced NH3 vapor outside the NH3 ice line in a fiducial model.
@@ 2.5 Planetary nitrogen abundance
We calculate the planetary nitrogen abundance sequentially from the disk evolution. For computational simplicity, we insert a protoplanet at time $t=t_{\rm p}$ and orbital radius $r=r_{\rm p}$, reset the inner boundary of the gas disk to the core’s orbit, and let the core accrete all the disk gas flowing to the inner computational boundary by applying the zero-torque boundary condition. The zero-torque boundary condition of the gas produces a pressure bump outside the inner boundary, which causes dust pile-up. This effect reproduces gap formation. We neglect core growth, planetary migration, and gas inflow from the inner region of the core’s orbit. This treatment is based on the runaway accretion scheme of Tanigawa & Tanaka (2016); Tanaka et al. (2020). They show analytically that a planet core with a 30 $M_{\oplus}$$10M_{\rm jup}$ would capture almost all gas inflowing into the planetary gap, where $M_{\rm jup}$ represents Jupiter’s mass. In our calculations, the protoplanet is assumed to have the initial mass $M_{\rm p,ini}$ of 30 $M_{\oplus}$ and comprises a heavy element core and an envelope. The mass of the heavy-element core is taken to be 13 $M_{\oplus}$ (0.04 $M_{\rm jup}$) based on Juno’s constraint that sets an upper limit of 6–27 $M_{\oplus}$ for the total heavy-element mass of Jupiter (Wahl et al., 2017; Ni, 2019). However, Jupiter’s internal structure model based on the Juno’s gravity data suggests that Jupiter’s core is diluted (e.g., Vazan et al., 2018; Debras & Chabrier, 2019; Helled et al., 2022). Furthermore, Jupiter’s formation models show that once the core grows into a few $M_{\oplus}$ in mass and acquires an envelop, heavy elements tend to dissolve in the envelope, irrespective of whether the solids are brought about by pebble accretion or planetesimal accretion (Iaroslavitz & Podolak, 2007; Hori & Ikoma, 2011; Brouwers et al., 2018; Valletta & Helled, 2020). Therefore, we assume that 10 $M_{\oplus}$ out of 13 $M_{\oplus}$ of heavy elements in the core dissolve into the atmosphere. Several long-term evolutionary models of Jupiter’s interior structure have demonstrated that convective mixing erases compositional gradients in the outer regions of the planet (50% or more of the radius) on scales of millions of years (Vazan et al., 2018; Müller et al., 2020). We assume that heavy elements dissolved in the atmosphere undergo homogenization through convective mixing. On the other hand, we note that the composition of actual Jupiter is not entirely uniform in the vertical direction, as evidenced by Jupiter having a diluted core. A discussion of compositional heterogeneity is provided in Section 4.6.
@@
We continue calculating the time evolution of dust and NH3 vapor using equations (7), (13) and (19), but with the inner boundary condition at $r=r_{\rm p}$. If the ammonium salt line is located inside the planetary core orbit, salt dissociation does not contribute to NH3 vapor abundance after core formation. For the dust flux at the inner boundary, we adopt a zero-flux condition because of the lack of background gas inside the boundary, i.e., no radial drift. The inner boundary flux of NH3 vapor is determined solely from the advection term: the first term on the right-hand side of equation (20).
@@ 2.6 Numerical settings and parameter set
The simulation consists of several models with different parameter sets. We set up as a fiducial model when a core forms at 0.6 Myr at 3.0 au in a disk with moderate turbulence strength $\alpha=10^{-3}$, a fragmentation velocity of icy dust $v_{\rm frag}=10$ m s-1, and salt dissociation temperature $T_{\rm vap,NH_{4}X}=200$ corresponding to ${\rm X}={\rm HCO_{2}}$. Other parameter sets are shown in Table 1, with seven cases in which the protoplanet formation age, core orbit, salt dissociation temperature, turbulence strength, or dust fragmentation velocity differ from the fiducial model. In addition to these parameter sets, we calculate Case 8, which considers the effect of recondensation according to equations (22) and (23). For each case, we take the ammonium salt content in the dust to be 10, 20, or 30 wt(corresponding to $f_{\rm NH_{4}X,dust}=0.2^{+0.1}_{-0.1}$). In all cases, the simulations run until the planet reaches Jupiter’s mass.
@@ 3.1.1 Time evolution of the disk until the core formation
Figure 2: Time evolution of the disk temperature $T$ from the fiducial model as a function of the distance from the central star $r$. The blue dotted, dashed, and solid lines are the snapshots at times $t=0,0.1$ and 0.6 Myr, respectively. The gray and green dash-dotted lines represent the sublimation temperature of NH3 and the dissociation temperature of NH4HCO2, respectively. Alt text: Line graph showing the disk temperature evolution.
@@
Figure 3: Time evolution of the gas surface density $\Sigma_{\rm g}$ (upper left), the dust surface density $\Sigma_{\rm d}$ (upper right), the Stokes number St (lower left), and the dust grain size (lower right), respectively. The results are from the fiducial model. The blue dotted, dashed, and solid lines are the snapshots at times $t=0,0.1$ and 0.6 Myr, respectively. The small drops in $\Sigma_{\rm d}$ at 2.5 au (on the dashed line) and 1.8 au (on the solid line) are caused by ammonium salt dissociation. Alt text: Four line graphs showing the evolution of physical properties of gas and dust.
@@
Figure 4: Time evolution of the number of NH3 per unit $\ln r$ derived from NH4HCO2 salt $N_{\rm NH_{4}HCO_{2},vap}^{\prime}$ (green) and NH3 ice $N_{\rm NH_{3},vap}^{\prime}$ (blue) from the fiducial model. The dashed and solid lines are the snapshots at $t=0.1$ and 0.6 Myr. The open and filled circles above the panel indicate the locations of the NH3 ice line (gray) and the ammonium salt line (black) at $t=0$ and $t=0.6$ Myr. Alt text: Line graph showing the distribution of NH3 molecules produced by dissociation or sublimation.
@@ 3.1.2 Gas accretion onto the protoplanet
Figure 5: Time evolution of the gas surface density $\Sigma_{\rm g}$ (left) and $N_{i,\rm vap}^{\prime}$ (right) after the core insertion. The core is inserted at $t_{\rm p}=$ 0.6 Myr and $r_{\rm p}=$ 3.0 au. These results are from the fiducial model. The dotted, dashed, and solid lines are the snapshots at times $t-t_{\rm p}=0,0.01$ and 0.04 Myr ($M_{\rm p}=M_{\rm jup}$), respectively. Here, $M_{\rm p}$ is the total planetary mass as $M_{\rm p}=M_{\rm p,atm}+M_{\rm p,core}$. Alt text: Two line graphs showing the time evolution of the gas surface density and the number of NH3 molecules after the core insertion.
@@
Figure 6: Mass of the planet as a function of time after core insertion $t-t_{\rm p}$ (left) and atmospheric nitrogen abundance of the planet $x_{\rm{N,p}}$ normalized by solar value of $7.78\times 10^{-5}$ (Asplund et al., 2021) as a function of $M_{\rm p}$ (right) from the fiducial model. The orange line in the right panel shows the planetary nitrogen abundance $x_{\rm N,p}$, with the orange shaded area representing the 10 wtuncertainty in the NH4HCO2 salt content in comets (Poch et al., 2020). The blue line indicates the contribution of accreted NH3 vapor only ($n_{\rm NH_{3},atm}/(n_{\rm H,atm}+3n_{\rm NH_{3},atm})$). The black hatched bar marks Jupiter’s atmospheric nitrogen abundance including observational errors from Juno (Bolton et al., 2017). The dot-dashed line marks Jupiter’s nitrogen abundance from Galileo probe (Wong et al., 2004); this value is highly uncertain ($4.2^{+1.6}_{-1.6}$), but the uncertainty is not indicated in the figure for clarity. The light gray line marks the lower limit of Saturn’s nitrogen abundance from Cassini (Fletcher et al., 2009). Alt text: Two line graphs showing the evolution of planetary mass and atmospheric nitrogen abundance.
@@
Here we describe disk evolution after core insertion and gas accretion by the core. Figure 5 shows the time evolution of $\Sigma_{\rm g}$ and $N_{i,\rm vap}^{\prime}$ as a function of the time $t-t_{\rm p}$ after core insertion. The NH3 ice and ammonium salt lines are located outside and inside the core orbit, respectively. Although there is a small supply of NH3 vapor after the core insertion, the amount of nitrogen vapor that accretes on the planet is largely determined by the amount of vapor that has reached the outer core orbit prior to the core insertion. The left panel of Figure 6 displays the evolution of the planet mass. The planet mass reaches one Jupiter mass $M_{\rm jup}$ within 0.04 Myr after the onset of gas accretion. This timescale is shorter than the growth time of Jupiter via runaway gas accretion estimated from numerical hydrodynamic simulations, but the difference is limited to a factor of a few, up to an order of magnitude. (Lissauer et al., 2009; Tanigawa & Tanaka, 2016; Lozovsky et al., 2017; Guillot et al., 2023).
@@
The right panel of Figure 6 shows the nitrogen enrichment of the planet $x_{\rm N,p}$ as a function of $M_{\rm p,atm}$. Here we plot the possible range of nitrogen abundance that was originally derived from the dissociation of the ammonium salt, sublimation of NH3 ice, and ammonium salt in the core, taking into account the 10 wtuncertainty in the abundance of salts contained in the disk dust. To highlight the contribution of NH3 vapor to nitrogen enrichment, we also plot the nitrogen abundance of accreting gas only. The final nitrogen abundance in the planetary atmosphere exceeds the solar abundance by a factor of 1–3. Dust containing 20 wtof ammonium salts yields $x_{\rm N,p}$ comparable to that measured by Juno for Jupiter (Bolton et al., 2017). Salts incorporated into the core and accreted NH3 vapor almost equally contribute to this degree of nitrogen enrichment; either one alone results in the enrichment degree below Juno’s observations. The nitrogen abundance decreases as the planet grows. This occurs because the contribution of core dissolution to nitrogen enrichment is already accounted for at the time of the protoplanet insertion, but it should be noted that this result does not constrain the timing of core dissolution during Jupiter’s growth.